# Minimal Construction of Graphs with Maximum Robustness

### v1

@@ 2.4 Problem Statement
In this paper, we aim to construct graphs with a given number of nodes that achieve the maximum $r$- and $(r,s)$- robustness while minimizing the number of edges. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $|\mathcal{V}|=n$. Before formally defining our problems, we first define two classes of graphs below:
Definition 9 ($\mathbf{r}$-MERG)
@@
Definition 10 ($\mathbf{(r,s)}$-MERG)
@@
The $r$-MERGs and $(r,s)$-MERGs are general classes of graphs that achieve certain robustness levels with minimal sets of edges. Note that in [37] we have provided similar definitions but due to the conflict of names, we are redefining with different names. In our paper, we are interested in finding the special classes of these, i.e., those that achieve maximum robustness with the fewest edges. Maximum $r$-robustness $\mathcal{G}$ can achieve is $r=\lceil\frac{n}{2}\rceil$ [4]. Using this fact, we now formally introduce two problems of our interest:
Problem 1
@@
Remark 1
While the actual maximum $(r,s)$-robustness $\mathcal{G}$ can achieve is $(\lceil\frac{n}{2}\rceil,n)$, this level can only be achieved by complete graphs [4, Lemma 4], which are rather trivial. Instead, we focus our analysis on the case where $r=s=\lceil\frac{n}{2}\rceil$, as $(\lceil\frac{n}{2}\rceil,\lceil\frac{n}{2}\rceil)$-robust network can tolerate up to $\lceil\frac{n}{2}\rceil-1$ malicious agents, which is the maximum number of malicious agents the network with $n$ nodes can tolerate through the W-MSR in practice.
For ease of notation, we define $\gamma=\lceil\frac{n}{2}\rceil$ for the remainder of the paper. It is important to note that analyzing graph structures with maximum robustness requires distinct analyses for even and odd values of $n$. In fact, we can describe $n$ in terms of $\gamma$ as follows: $n=2\gamma-1$ for odd $n$, and $n=2\gamma$ for even $n$. Note that for a fixed $\gamma$, the even case yields an $n$ value one greater than the odd case. Despite this distinction, the solution procedures remain the same. Our analysis takes place in two main steps: First, we establish the lower bounds on the edge counts for a graph with $n$ nodes to achieve maximum robustness ($\gamma$- and $(\gamma,\gamma)$-robustness). Second, we leverage these lower bounds to develop construction mechanisms that are proven to construct $\gamma$- and $(\gamma,\gamma)$-MERGs.
@@ 3 Bounds on Edge Counts for $\gamma$-MERGs
This section addresses the first step toward constructing $\gamma$-MERGs by examining the minimum number of edges required for a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=n$ to achieve $\gamma=\lceil\frac{n}{2}\rceil$-robustness. Due to the ceiling function involving $\gamma$, it is necessary to separately analyze cases where $n$ is odd or even. Thus, this section is divided into two cases. For the first case, we find the lower bounds on edge counts for $\gamma$-robust graphs when $n$ is odd. Then, for the second case, we do the same for $\gamma$-robust graphs when $n$ is even. These are formally established in Lemma 3.2 and Lemma 3.10, respectively. These two lemmas will be later used in Section 4 to construct $\gamma$-MERGs with any $n$.
@@
In general, $k$-clique refers to a clique consisting of $k$ nodes. The clique with the largest cardinality in $\mathcal{G}$ is called the maximum clique. Cliques are the building blocks of our analysis in this section, as we examine the necessary structures of $r$-robust graphs in general graph settings. We will use cliques to discover some necessary graph structures within $\gamma$-robust graphs, which will help us establish the lower bounds on edge counts.
@@ 3.1 Case 1: Graphs with Odd $n$
We now examine the minimum required size of the largest clique that any $\gamma$-robust graph with an odd number of nodes $n$ must contain in order to satisfy the conditions of $r$-robustness (Definition 6).
@@
Proof 3.1.
In this proof, we will show our argument by constructing two subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that $\mathcal{G}$ contains a $(\gamma+1)$-clique. Since $\mathcal{G}$ is $\gamma$-robust, it holds that for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ of $\mathcal{V}$, at least one of them is $\gamma$-reachable. Hence, by Definition 6, for any node in $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, $|\mathcal{N}_{v_{i}}|\geq\gamma$ must hold. WLOG, let $v_{1}\in\mathcal{V}$ be a node, as shown on Figure 1 (a). Then because $n=2\gamma-1$, we can construct $\mathcal{S}_{1}$ to be the set of $\gamma-1$ neighbors of $v_{1}$ and $\mathcal{S}_{2}$ to be the set of the remaining $\gamma$ nodes including $v_{1}$ such that $|\mathcal{S}_{1}|=\gamma-1$, $|\mathcal{S}_{2}|=\gamma$, and $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$. This enforces $\mathcal{S}_{1}$ to be $\gamma$-reachable.
@@
Likewise, in the $k^{\text{th}}$ step, $\exists v_{k+1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{k+1}}\cap\mathcal{S}_{2}|\geq\gamma$ and thus forms a $(k+1)$-clique with $v_{1},v_{2},\cdots,v_{k-1},v_{k}\in\mathcal{S}_{2}$. We then swap $\gamma-k$ nodes in $\mathcal{S}_{1}$ including $v_{k+1}$ with $\gamma-k$ nodes in $\mathcal{S}_{2}\setminus\{v_{1},\cdots,v_{k}\}$. At the end of $\gamma-1^{\text{th}}$ step, $\mathcal{S}_{2}$ contains $v_{1},\cdots,v_{\gamma}$ that form a $\gamma$-clique. Finally, since $\mathcal{S}_{1}$ is $\gamma$-reachable, it contains a node $v_{\gamma+1}$ that has edges with all nodes $v_{1},\cdots,v_{\gamma}\in\mathcal{S}_{2}$. Then, $\{v_{\gamma+1}\}\cup\mathcal{S}_{2}$ forms a ($\gamma+1)$-clique.
Lemma 1 states that for a graph with odd $n$ to achieve $\gamma$-robustness (maximum $r$-robustness), it must contain a clique size of $\gamma+1$ or more nodes. This effectively builds a necessary subgraph structure for further analysis. Note that the existence of $(\gamma+1)$-clique implies that $\mathcal{G}$ must have as many edges as to construct a $(\gamma+1)$-clique, i.e., $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}$. Stemming from this reasoning, we finalize the lower bound on edge counts for the case of odd $n$:
Lemma 3.2.
@@
Proof 3.3.
Since $\mathcal{G}$ is $\gamma$-robust, it holds that for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}\subset\mathcal{V}$, at least one of them is $\gamma$-reachable (by Definition 6). From Lemma 1, we know that $\mathcal{V}$ must contain a $(\gamma+1)$-clique $\mathcal{C}$. Let $\mathcal{C}^{\prime}=\mathcal{V}\setminus\mathcal{C}$. This then implies $|\mathcal{C}|=\gamma+1$ and $|\mathcal{C}^{\prime}|=\gamma-2$.
WLOG, let $\mathcal{S}_{1}=\mathcal{C}^{\prime}$ and $\mathcal{S}_{2}=\mathcal{C}$. Then, $|\mathcal{S}_{1}|=\gamma-2$ and $|\mathcal{S}_{2}|=\gamma+1$. This implies $\mathcal{S}_{1}$ has to be $\gamma$-reachable. In other words, there must exist $v_{1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{1}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{v_{1}}\setminus\mathcal{S}_{1}|\geq\gamma$. Therefore, for such $v_{1}$ to exist, we need $\gamma$ edges that connect $v_{1}$ with $\gamma$ nodes in $\mathcal{S}_{2}$: $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}+\gamma$. Note that $v_{1}\in\mathcal{S}_{1}$ is the node that makes $\mathcal{S}_{1}$ $\gamma$-reachable. Therefore, moving $v_{1}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$ requires the new $\mathcal{S}_{1}$, which is now without $v_{1}$, to contain another node $v_{2}$ that has at least $\gamma$ neighbors in $\mathcal{S}_{2}$. This again requires at least $\gamma$ additional edges (i.e. $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}+2\gamma$). We then move $v_{2}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$, which necessitates $\mathcal{S}_{1}$ to have another node $v_{3}$ to have $\gamma$ neighbors in $\mathcal{S}_{2}$.
We continue this process of (i) connecting a node $v_{i}\in\mathcal{S}_{1}$ with any $\gamma$ nodes in $\mathcal{S}_{2}$ and (ii) moving $v_{i}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$ in the $i^{\text{th}}$ step until $\mathcal{S}_{1}$ becomes empty. Then, we can repeat these two steps $\gamma-2$ times, as we initially had $|\mathcal{S}_{1}|=\gamma-2$. In other words, for each $\gamma-2$ pair of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, we require $\gamma$ additional edges. Since we started from the setting where $\mathcal{S}_{2}=\mathcal{C}$ and forced $|\mathcal{E}|$ to increase at least a total of $\gamma(\gamma-2)$, $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)+2\gamma(\gamma-2)}{2}\Rightarrow|\mathcal{E}|\geq\frac{3\gamma(\gamma-1)}{2}$.
From Lemma 3.2, we can conclude that any $\gamma$-robust graphs with odd $n$, i.e., graphs with an odd number of nodes that achieve the maximum $r$-robustness, must have $\frac{3\gamma(\gamma-1)}{2}$ or more edges. In fact, Theorem 4.16 later shows that the bound in the lemma is tight and cannot be further improved. Furthermore, using Lemma 3.2, we can derive a useful corollary:
Corollary 3.4.
@@
Proof 3.5.
In general, the more nodes a graph has, the more edges it must have to maintain a given robustness. From [30], it is shown that $12r-$ 1 is the smallest number of nodes a graph must have to be $r$-robust. In other words, $r$-robustness is the maximum robustness a graph with $12r-$ 1 nodes can achieve. Thus, Lemma 3.2 can be restated: any $r$-robust graph with $12r-$ 1 nodes must have at least $\frac{3r(r-1)}{2}$ edges. Consequently, any graph $\mathcal{G}$ with $12r-$ 1 or more nodes must have at least $\frac{3r(r-1)}{2}$ edges to be $r$-robust.
Remark 3.6.
Corollary 3.4 reveals that the bound given in Lemma 3.2 can be applied beyond $\gamma$-robustness. The bound in fact serves as a necessary condition for any $r$-robust graphs of any number of nodes $n$, ensuring that graphs with edges less than $\frac{3r(r-1)}{2}$ cannot be $r$-robust. One useful implication of this result is that it provides a quick and easy way to upper-bound a graph’s robustness level based solely on its number of edges. Computing $r$-robustness of a graph is in general a very difficult problem [8, 21, 42]. While various efficient methods have been proposed to approximate robustness levels [22, 23, 24], these approaches would still involve significant computational effort as graph sizes increase and frequently lack formal guarantees. In contrast, verifying whether a given graph satisfies the bound (2) is computationally efficient. It can also accelerate existing methods by reducing their search space through the upper bound on robustness provided by this result beforehand.
@@ 3.2 Case 2: Graphs with Even $n$
Now we continue our analysis on the edge counts for the graphs with even $n$. Since $n$ is even, $\gamma=\frac{n}{2}$. This implies that, for the same value of $\gamma$, the value of $n$ is one greater when $n$ is even compared to when it is odd. Because there is one more node, the analysis involves more steps. However, the overall approach is similar to the previous case: we will first determine the minimum size of the maximum clique that $\mathcal{G}$ must contain to be $\gamma$-robust.
Lemma 3.7.
@@
Due to length, the proof for Lemma 3.7 is given in the Appendix (Section 9). Comparing Lemma 1 and 3.7, it can be seen that the size of the necessary maximum clique in a $\gamma$-robust graph we can guarantee decreases as its number of node increases. This is reasonable, as the number of possible graph configurations increases exponentially with the number of nodes, thereby making the analysis more complex. Due to the increased complexity, unlike the analysis in the previous subsection for the odd $n$ case, the study of $\gamma$-robust graphs with even $n$ requires topological structures beyond the maximum clique. The following lemma presents the additional necessary structures.
Lemma 3.8.
@@
Proof 3.9.
Since $\mathcal{G}$ is $\gamma$-robust, for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}\subset\mathcal{V}$, at least one of them is $\gamma$-reachable. From Lemma 3.7, $\mathcal{G}$ contains a $\big{(}\lfloor\frac{\gamma+4}{2}\rfloor\big{)}$-clique $\mathcal{C}$. WLOG, let $\mathcal{S}_{2}$ contain $\gamma+1$ nodes including $\mathcal{C}$, and let $\mathcal{S}_{1}$ contain the remaining $\gamma-1$ nodes. Initially, we set $\mathcal{V}_{S}=\mathcal{C}$. Then, $|\mathcal{E}_{S}|=\big{(}\frac{1}{2}\big{)}\big{(}\lfloor\frac{\gamma+4}{2}\rfloor\big{)}\big{(}\lfloor\frac{\gamma+2}{2}\rfloor\big{)}$. Note there are total $n=2\gamma$ nodes. Since $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$, $\mathcal{S}_{1}$ must be $\gamma$-reachable.
By Lemma 3.7, we have a $2$-clique and $3$-clique when $\gamma=1$ and $\gamma=2$ respectively, satisfying the statement in the lemma for $\gamma=1,2$.
@@
We remove the floor functions: if $\gamma$ is even, $|\mathcal{E}_{S}|\geq\big{(}\frac{1}{2}\big{)}\big{(}\frac{\gamma+4}{2}\big{)}\big{(}\frac{\gamma+2}{2}\big{)}+\big{(}\frac{\gamma+2}{2}\big{)}+\big{(}\frac{\gamma+4}{2}\big{)}+\cdots+(\gamma-2)+(\gamma-1)=\frac{\gamma^{2}+2}{2}$. If $\gamma$ is odd, $|\mathcal{E}_{S}|\geq\big{(}\frac{1}{2}\big{)}\big{(}\frac{\gamma+3}{2}\big{)}\big{(}\frac{\gamma+1}{2}\big{)}+\big{(}\frac{\gamma+1}{2}\big{)}+\big{(}\frac{\gamma+3}{2}\big{)}+\cdots+(\gamma-2)+(\gamma-1)=\frac{\gamma^{2}+1}{2}$. Thus, we have $|\mathcal{V}_{S}|=\gamma+1$ and $|\mathcal{E}_{S}|\geq\lfloor{\frac{\gamma^{2}+2}{2}}\rfloor$.
Lemma 3.8 effectively provides a necessary induced subgraph structure that $\mathcal{G}$ with even $n$ must have to be $\gamma$-robust. Now, using this lemma, we present the following:
Lemma 3.10.
@@
Proof 3.11.
The approach of the proof is identical to that of Lemma 3.2, except for the fact that in this proof, we are starting with an induced subgraph $\mathcal{G}_{S}=(\mathcal{V}_{S},\mathcal{E}_{S})$ such that $|\mathcal{V}_{S}|=\gamma+1$ and $|\mathcal{E}_{S}|\geq\lfloor{\frac{r^{2}+2}{2}}\rfloor$ (which comes from Lemma 3.8).

### v2

@@ 2.3 Problem Statement
In this paper, we aim to construct graphs with a given number of nodes that achieve maximum $r$- and $(r,s)$- robustness using the minimal number of edges. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $|\mathcal{V}|=n$. Before formally defining our problems, we first define two classes of graphs below:
Definition 9 ($\mathbf{r}$-MERG).
@@
Definition 10 ($\mathbf{(r,s)}$-MERG).
@@
The $r$-MERGs and $(r,s)$-MERGs are general classes of graphs that achieve certain robustness levels with minimal sets of edges. Note that in [45] we have provided definitions of the same properties using different terms, which however may cause ambiguity and appear conflicting to the well-established notion of $(r,s)$-robust graphs. Hence in this paper we introduce the MERG definitions above to avoid any confusion. Since the maximum $r$-robustness $\mathcal{G}$ can achieve is $r=\lceil\frac{n}{2}\rceil$ [4], the problems addressed in this paper are:
Problem 1.
@@
Remark 1.
While the actual maximum $(r,s)$-robustness $\mathcal{G}$ achieves is $(\lceil\frac{n}{2}\rceil,n)$, this is only achieved by complete graphs [4, Lemma 4], which are rather trivial. Instead, we focus on the case where $r=s=\lceil\frac{n}{2}\rceil$, as $(\lceil\frac{n}{2}\rceil,\lceil\frac{n}{2}\rceil)$-robust network tolerates up to $\lceil\frac{n}{2}\rceil-1$ malicious agents, which is the maximum number of malicious agents the network with $n$ nodes tolerates through the W-MSR algorithm in practice.
For ease of notation, we define $\gamma=\lceil\frac{n}{2}\rceil$ for the remainder of the paper. Systematically constructing $\gamma$- and $(\gamma,\gamma)$-MERGs allows us to design networks that achieve the highest possible resilience for many distributed algorithms to misbehaving agents with the minimal number of communication links. Such constructions allow the system across diverse application domains to enjoy the highest robustness with reduced communication and energy overhead, which is critical for resource-constrained networks such as sensor networks, robotic swarms, and other distributed control systems. Moreover, the proposed constructions provide graph topologies that also work for extensions of the W-MSR algorithm which address practical network constraints, such as time-varying topologies and limited-rate communication [50, 39], as well as multi-dimensional states [51], without requiring additional edges.
Before proceeding with our analyses, we note that for both $\gamma$- and $(\gamma,\gamma)$-MERGs, we separate the analyses into two cases based on whether $n$ is even or odd. This separation is made for analytical convenience in our approach. For both cases, our analysis takes place in two steps: First, we establish the lower bounds on the edge counts for a graph with $n$ nodes to achieve maximum robustness ($\gamma$- and $(\gamma,\gamma)$-robustness). Second, we leverage these lower bounds to develop construction mechanisms that are proven to construct $\gamma$- and $(\gamma,\gamma)$-MERGs.
@@ 3 Bounds on Edge Counts for $\gamma$-MERGs
This section addresses the first step toward constructing $\gamma$-MERGs by examining the minimum number of edges required for a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=n$ to achieve $\gamma=\lceil\frac{n}{2}\rceil$-robustness. This section consider two cases. In the first case, we find the lower bounds on edge counts for $\gamma$-robust graphs with odd $n$. Then, in the second case, we do the same for $\gamma$-robust graphs with even $n$. These are formally established in Lemma 2 and Lemma 5, respectively. These lemmas will be later used to construct $\gamma$-MERGs in Section 4.
@@
In general, $k$-clique refers to a clique consisting of $k$ nodes. The clique with the largest cardinality in $\mathcal{G}$ is called the maximum clique. We will use cliques to discover necessary graph structures within $\gamma$-robust graphs, which will help us establish the lower bounds on edge counts.
@@ 3.1 Case 1: Graphs with Odd $n$
We now examine the lower bounds on the number of edges for the $\gamma$-robust graphs with odd $n$. To this end, we first study the smallest maximum clique that a $\gamma$-robust graph with odd $n$ must contain in order to satisfy the conditions of $r$-robustness (Definition 6).
@@
Proof.
In this proof, we will show our argument by constructing two subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that $\mathcal{G}$ contains a $(\gamma+1)$-clique. Since $\mathcal{G}$ is $\gamma$-robust, it holds that for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ of $\mathcal{V}$, at least one of them is $\gamma$-reachable. Hence, by Definition 6 and [4, Lemma 5], for any node in $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, $|\mathcal{N}_{v_{i}}|\geq\gamma$ must hold. WLOG, let $v_{1}\in\mathcal{V}$ be a node, as shown on Figure 1 (a). Then because $n=2\gamma-1$, we can construct $\mathcal{S}_{1}$ to be the set of $\gamma-1$ neighbors of $v_{1}$ and $\mathcal{S}_{2}$ to be the set of the remaining $\gamma$ nodes including $v_{1}$ such that $|\mathcal{S}_{1}|=\gamma-1$, $|\mathcal{S}_{2}|=\gamma$, and $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$. This enforces $\mathcal{S}_{1}$ to be $\gamma$-reachable.
@@
Likewise, in the $k^{\text{th}}$ step, $\exists v_{k+1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{k+1}}\cap\mathcal{S}_{2}|\geq\gamma$ and thus forms a $(k+1)$-clique with $v_{1},v_{2},\cdots,v_{k-1},v_{k}\in\mathcal{S}_{2}$. We then swap $\gamma-k$ nodes in $\mathcal{S}_{1}$ including $v_{k+1}$ with $\gamma-k$ nodes in $\mathcal{S}_{2}\setminus\{v_{1},\cdots,v_{k}\}$. At the end of $\gamma-1^{\text{th}}$ step, $\mathcal{S}_{2}$ contains $v_{1},\cdots,v_{\gamma}$ that form a $\gamma$-clique. Finally, since $\mathcal{S}_{1}$ is $\gamma$-reachable, it contains a node $v_{\gamma+1}$ that has edges with all nodes $v_{1},\cdots,v_{\gamma}\in\mathcal{S}_{2}$. Then, $\{v_{\gamma+1}\}\cup\mathcal{S}_{2}$ forms a ($\gamma+1)$-clique. ∎
Lemma 1 states that for a graph with odd $n$ to be $\gamma$-robust, it must contain a clique size of $\gamma+1$ or more nodes. Stemming from this reasoning, we finalize the lower bound on edge counts for the case of odd $n$:
Lemma 2.
@@
Proof.
Since $\mathcal{G}$ is $\gamma$-robust, it holds that for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}\subset\mathcal{V}$, at least one of them is $\gamma$-reachable (by Definition 6). From Lemma 1, we know that $\mathcal{V}$ must contain a $(\gamma+1)$-clique $\mathcal{C}$. This means that $\mathcal{G}$ must have at least as many edges as to construct a $(\gamma+1)$-clique, i.e., $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}$.
Let $\mathcal{C}^{\prime}=\mathcal{V}\setminus\mathcal{C}$. This then implies $|\mathcal{C}|=\gamma+1$ and $|\mathcal{C}^{\prime}|=\gamma-2$. WLOG, let $\mathcal{S}_{1}=\mathcal{C}^{\prime}$ and $\mathcal{S}_{2}=\mathcal{C}$. Then, $|\mathcal{S}_{1}|=\gamma-2$ and $|\mathcal{S}_{2}|=\gamma+1$. This implies $\mathcal{S}_{1}$ has to be $\gamma$-reachable. In other words, there must exist $v_{1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{v_{1}}\cap\mathcal{S}_{2}|=|\mathcal{N}_{v_{1}}\setminus\mathcal{S}_{1}|\geq\gamma$. Therefore, for such $v_{1}$ to exist, we need $\gamma$ edges that connect $v_{1}$ with $\gamma$ nodes in $\mathcal{S}_{2}$: $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}+\gamma$. Note that $v_{1}\in\mathcal{S}_{1}$ is the node that makes $\mathcal{S}_{1}$ $\gamma$-reachable. Therefore, moving $v_{1}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$ requires the new $\mathcal{S}_{1}$, which is now without $v_{1}$, to contain another node $v_{2}$ that has at least $\gamma$ neighbors in $\mathcal{S}_{2}$. This again requires at least $\gamma$ additional edges (i.e. $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)}{2}+2\gamma$). We then move $v_{2}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$, which necessitates $\mathcal{S}_{1}$ to have another node $v_{3}$ to have $\gamma$ neighbors in $\mathcal{S}_{2}$.
We continue this process of (i) connecting a node $v_{i}\in\mathcal{S}_{1}$ with any $\gamma$ nodes in $\mathcal{S}_{2}$ and (ii) moving $v_{i}$ from $\mathcal{S}_{1}$ to $\mathcal{S}_{2}$ in the $i^{\text{th}}$ step until $\mathcal{S}_{1}$ becomes empty. Then, it is possible to repeat these two steps $\gamma-2$ times, as we initially had $|\mathcal{S}_{1}|=\gamma-2$. In other words, for each $\gamma-2$ pair of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, we require $\gamma$ additional edges. Since we started from the setting where $\mathcal{S}_{2}=\mathcal{C}$ and forced $|\mathcal{E}|$ to increase at least a total of $\gamma(\gamma-2)$, $|\mathcal{E}|\geq\frac{\gamma(\gamma+1)+2\gamma(\gamma-2)}{2}\Rightarrow|\mathcal{E}|\geq\frac{3\gamma(\gamma-1)}{2}$. ∎
Using Lemma 2, we conclude that any $\gamma$-robust graphs with odd $n$, i.e., graphs with an odd number of nodes that achieve maximum $r$-robustness, must have $\frac{3\gamma(\gamma-1)}{2}$ or more edges. In fact, Theorem 1 later shows that the bound in the lemma is tight and cannot be further improved. Furthermore, using Lemma 2, we derive a useful corollary:
Corollary 1.
@@
Proof.
In general, the more nodes a graph has, the more edges it must have to maintain a given robustness. From [37], it is shown that 2r12r-1 is the smallest number of nodes a graph must have to be $r$-robust. In other words, $r$-robustness is the maximum robustness a graph with 2r12r-1 nodes achieves. Thus, Lemma 2 is restated as: any $r$-robust graph with 2r12r-1 nodes must have at least $\frac{3r(r-1)}{2}$ edges. Consequently, any graph $\mathcal{G}$ with 2r12r-1 or more nodes must have at least $\frac{3r(r-1)}{2}$ edges to be $r$-robust. ∎
Remark 2.
Corollary 1 reveals that the bound in Lemma 2 is in fact a necessary condition for any $r$-robust graphs of any number of nodes $n$, ensuring that graphs with edges less than $\frac{3r(r-1)}{2}$ cannot be $r$-robust. One useful implication of this result is that it provides a quick way to upper-bound a graph’s robustness level based solely on its number of edges. Directly computing $r$-robustness of a graph is generally difficult [8, 24], and while various approximation methods exist [25, 26, 27], they often lack formal guarantees or assume some structures on the given graphs. In fact, the upper bound on robustness derived from this result can be used to accelerate existing methods by reducing their search space.
@@ 3.2 Case 2: Graphs with Even $n$
Now we continue our analysis on the edge counts for the graphs with even $n$. Since $n$ is even, $\gamma=\frac{n}{2}$. This implies that, for the same value of $\gamma$, the value of $n$ is one greater when $n$ is even compared to when it is odd. However, the overall approach is similar to the previous case: we will first determine the minimum size of the maximum clique that $\mathcal{G}$ must contain to be $\gamma$-robust.
Lemma 3.
@@
Due to length, the proof for Lemma 3 is given in the Appendix (Section 10). We emphasize that the bounds from Lemmas 1 and 3 are only analytical lower bounds we are able to show, and may not reflect the actual necessary maximum clique size. In fact, unlike the analysis for odd $n$, studying $\gamma$-robust graphs with even $n$ involves topological structures beyond the maximum clique we observe in Lemma 3. The following lemma presents the additional necessary structures.
Lemma 4.
@@
Proof.
Since $\mathcal{G}$ is $\gamma$-robust, for any pair of disjoint, non-empty subsets $\mathcal{S}_{1}$, $\mathcal{S}_{2}\subset\mathcal{V}$, at least one of them is $\gamma$-reachable. From Lemma 3, $\mathcal{G}$ contains a $\big(\lfloor\frac{\gamma+4}{2}\rfloor\big)$-clique $\mathcal{C}$. WLOG, let $\mathcal{S}_{2}$ contain $\gamma+1$ nodes including $\mathcal{C}$, and let $\mathcal{S}_{1}$ contain the remaining $\gamma-1$ nodes. Initially, we set $\mathcal{V}_{S}=\mathcal{C}$. Then, $|\mathcal{E}_{S}|=\big(\frac{1}{2}\big)\big(\lfloor\frac{\gamma+4}{2}\rfloor\big)\big(\lfloor\frac{\gamma+2}{2}\rfloor\big)$. Note there are total $n=2\gamma$ nodes. Since $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$, $\mathcal{S}_{1}$ must be $\gamma$-reachable.
By Lemma 3, we have a $2$-clique and $3$-clique when $\gamma=1$ and $\gamma=2$ respectively, satisfying the statement in the lemma for $\gamma=1,2$.
@@
We remove the floor functions: if $\gamma$ is even, $|\mathcal{E}_{S}|\geq\big(\frac{1}{2}\big)\big(\frac{\gamma+4}{2}\big)\big(\frac{\gamma+2}{2}\big)+\big(\frac{\gamma+2}{2}\big)+\big(\frac{\gamma+4}{2}\big)+\cdots+(\gamma-2)+(\gamma-1)=\frac{\gamma^{2}+2}{2}$. If $\gamma$ is odd, $|\mathcal{E}_{S}|\geq\big(\frac{1}{2}\big)\big(\frac{\gamma+3}{2}\big)\big(\frac{\gamma+1}{2}\big)+\big(\frac{\gamma+1}{2}\big)+\big(\frac{\gamma+3}{2}\big)+\cdots+(\gamma-2)+(\gamma-1)=\frac{\gamma^{2}+1}{2}$. Thus, we have $|\mathcal{V}_{S}|=\gamma+1$ and $|\mathcal{E}_{S}|\geq\lfloor{\frac{\gamma^{2}+2}{2}}\rfloor$. ∎
Lemma 4 effectively provides a necessary induced subgraph structure that $\mathcal{G}$ with even $n$ must have to be $\gamma$-robust. Now, using this lemma, we present the following:
Lemma 5.
@@
Proof.
The approach of the proof is identical to that of Lemma 2, except for the fact that in this proof, we are starting with an induced subgraph $\mathcal{G}_{S}=(\mathcal{V}_{S},\mathcal{E}_{S})$ such that $|\mathcal{V}_{S}|=\gamma+1$ and $|\mathcal{E}_{S}|\geq\lfloor{\frac{r^{2}+2}{2}}\rfloor$ (which comes from Lemma 4).